in a class of 68 students, there are 8 fewer democrats than republicans, and there are only half as many independents as democrats if every student is registered as a democrat, republican, or independent, how many students are in each group?
Answer by checkley77(12844)

R+(R-8)+(R-8)/2=68
R+R-8+.5R-4=68
2.5R=68+12
2.5R=80
R=80/2.5
R=32 REPUBLICANS.
32-8=24 DEMOCRATS.
24.2=12 INDEPENDENTS.
PROOF:
32+24+12=68
68=68